OpenStudy (anonymous):

Hey wassup homies i just had a question for u guys today, so lets do this! Which is greater: 0.001 or 0.01

OpenStudy (anonymous):

FAN AND MEDAL

OpenStudy (anonymous):

.01

OpenStudy (anonymous):

.01=10/100 .001=1/100

OpenStudy (anonymous):

good job homeboy u just earned urself a fan and a medal

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome